OpenStudy (anonymous):

Solve the system of equations. 5x + y = 9 3x + 2y = 4 (-2, 5) (1, 4) (2, -1) (4, -4)

OpenStudy (luigi0210):

You could solve by substitution

OpenStudy (texaschic101):

easiest way would be to isolate y in the first equation then sub that back into the 2nd equation. Thus, solving by substitution

OpenStudy (texaschic101):

lets me help you get started... 5x + y = 9 -- subtract 5x from both sides y = -5x + 9 now sub -5x + 9 in for y in the 2nd equation and solve for x. Now, show me what you can do :)

OpenStudy (anonymous):

@sourwing

OpenStudy (luigi0210):

If you like elimination: \[\LARGE 5x + y = 9\] \[\LARGE 3x + 2y = 4\] *-2 \[\LARGE -10x-2y=-18\] \[\LARGE 3x+2y=4\] And staph it, stop being lazy and do some work c:

OpenStudy (anonymous):

dont feel like doing it im sick in the hospital so no i dont feel the need to

OpenStudy (doc.brown):

That's the first wrong answer I've seen on openstudy.

OpenStudy (anonymous):

lol^

OpenStudy (anonymous):

then whats the correct answer

OpenStudy (texaschic101):

I am sick too....so I guess I don't need to give you the answer

OpenStudy (doc.brown):

Last time I was in the hospital I taught myself linear algebra.

OpenStudy (doc.brown):

Which, by the way, says:\[\left[\begin{matrix}5 & 1 & 9 \\ 3 & 2 & 4\end{matrix}\right]\]\[\left[\begin{matrix}1 & \frac{1}{5} & \frac{1}{9} \\ 3 & 2 & 4\end{matrix}\right]\]\[\left[\begin{matrix}1 & \frac{1}{5} &\frac{1}{9} \\ 0 & \frac{7}{5} & \frac{-7}{5}\end{matrix}\right]\]\[\left[\begin{matrix}1 & 0&2 \\ 0 & 1&-1\end{matrix}\right]\]x=2 y=-1

OpenStudy (anonymous):

cool i dont care okay im in the HOSPITAL for a reason not at home!?!?!? DUHHHHH so i dont know why your complaining @texaschic101

OpenStudy (texaschic101):

I understand your in the hospital for a reason, and I am sorry about that. But that doesn't mean you are entitled to get all the answers and do none of the work. Thats all I am saying. And if you didn't feel you needed to do the work, then I didn't feel the need to give you the answer.

OpenStudy (anonymous):

Soo, whats the answer? @texaschic101 XD